OpenStudy (anonymous):

A fair coin is tossed 5 times in a row. The exact probability of the coin landing heads exactly 2 times is? #medal given

OpenStudy (anonymous):

@Error1603 @Kitten_is_back

OpenStudy (anonymous):

@Liz_Beauty45 can you help me pleasse

OpenStudy (liz_beauty45):

sure but just know i suck at math ;)

OpenStudy (anonymous):

ohhhh lol

OpenStudy (liz_beauty45):

so how can i help

OpenStudy (anonymous):

@purple_pink can you help me

OpenStudy (liz_beauty45):

oh okay wiat brb

OpenStudy (anonymous):

@jigglypuff314 can you help me

OpenStudy (liz_beauty45):

Hey does this help..The answer is 10/32=5/16. Here's how. At each step the choice is either heads or tails. First toss, H or T. Second toss, HH HT TH TT (example:first toss was H, second could be H or T and so on)

OpenStudy (purple_pink):

lol i dislike doing probabilities yet i am kinda good at it, lemme see how i can be of assistance.

OpenStudy (anonymous):

okay I will give you a medal @purple_pink

OpenStudy (liz_beauty45):

Did my answer help @klc097 if not im very sorry..

OpenStudy (liz_beauty45):

10/32/=5/16 makes since right

OpenStudy (liz_beauty45):

ppl theese days...gtg

OpenStudy (purple_pink):

i'm sorry i think i calculated wrong @jabez177

OpenStudy (anonymous):

ahh dang

OpenStudy (purple_pink):

maybe my brother could help you :/ he is doing "work samples"

OpenStudy (purple_pink):

@KendrickLamar2014 @skullpatrol

OpenStudy (purple_pink):

@rebeccaxhawaii @sweetburger @haleyelizabeth2017

OpenStudy (kendricklamar2014):

@Wendy.Ivette11714

OpenStudy (kendricklamar2014):

@jabez177

OpenStudy (anonymous):

im trying to pass this test I have 4 questions left after this one

OpenStudy (wendy.ivette11714):

Two heads can be in a toss of 5 times in 5C2 = 10 ways Each way(number of times at least two heads will come) has a probability of 1/32. Hence 10*1/32 = 5/16

OpenStudy (wendy.ivette11714):

Wait, we cant help on Test's @klc097

OpenStudy (anonymous):

its a PRE-test

OpenStudy (anonymous):

If the number of equally likely sample outcomes of a single stage of an experiment is 4, what is the total number of elements in the sample space if the experiment has 3 stages?

OpenStudy (wendy.ivette11714):

Oh well then its ok

OpenStudy (anonymous):

can you help me with this one? @Wendy.Ivette11714

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@Wendy.Ivette11714 fan me so I can message you

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (wendy.ivette11714):

I have to do an assigment rn.

jabez177 (jabez177):

Okay! :) Let me have a looksie! :-)

jabez177 (jabez177):

@Liz_Beauty45 is correct. :)